# A Survey of Reinforcement Learning For Economics

### 09 Mar 2026 v1

@@ 5.1 Setup
The environment is a $5\times 5$ deterministic gridworld with five actions (up, down, left, right, stay). Each non-terminal step incurs a penalty of 0.1-0.1; reaching the goal in the bottom-right corner yields +10+10. A small state-dependent perturbation ($\epsilon=0.001$) breaks the symmetry between equally-long paths, ensuring a unique optimal policy. [^77] Without the perturbation, any shortest path from start to goal is optimal. The perturbation subtracts $\epsilon\cdot i/N^{2}$ from each step reward, where $i$ is the destination state index. Since moving down increases the state index by $N$ while moving right increases it by 1, the agent strictly prefers right over down at every interior state. The discount factor is $\gamma=0.95$ and each episode is capped at 50 steps. The optimal path requires 8 steps, yielding an optimal discounted return of 9.30.787878Shared hyperparameters across all learning algorithms: learning rate $\alpha=0.1$ (decay 0.99995/episode) for value-based methods, $\alpha=0.01$ for policy gradient methods; $\epsilon$-greedy exploration from 1.0 to 0.01 (decay 0.9995/episode); trace decay $\lambda=0.9$; NPG actor step size $\eta=0.2$, critic step size 0.5 (both decay 0.99995/episode). DQN uses a two-layer network with 64 hidden units, replay buffer of 50,000 transitions, batch size 64, learning rate $10^{-3}$, target network update every 100 episodes, and $\epsilon$-decay 0.9999. PPO uses clip ratio 0.2, 4 optimization epochs per batch, GAE $\lambda=0.95$, learning rate $0.01$ (decay 0.99995/episode). All learning algorithms run for 500,000 episodes with seed 42.
@@ 5.2 Planning Versus Learning
VI converges in 9 iterations (0.001 seconds). On this deterministic gridworld, the iteration count matches the MDP diameter: the longest shortest path from any state to the goal is 8 steps, and each Bellman backup propagates value one step outward from the terminal state. [^79] The $\gamma^{k}$ contraction bound from Section 4.1 predicts roughly $\log(\epsilon^{-1})/\log(\gamma^{-1})\approx 360$ iterations for $\epsilon=10^{-8}$, making it loose by a factor of 40 on this structured problem.
@@ 5.3 Value Function Convergence
Q-learning and Q($\lambda$) are off-policy: they learn $Q^{*}$ regardless of the behavior policy. Both achieve $\max_{s}|V(s)-V^{*}(s)|<0.1$ by episode 610 and 550 respectively, with final RMSE below 0.01. Q($\lambda$) converges slightly faster due to eligibility traces propagating the terminal reward backward through the full 8-step trajectory within a single episode.
@@
SARSA, REINFORCE, NPG, and PPO are on-policy: they learn the value of the exploration policy $\pi_{\epsilon}$, not $\pi^{*}$. SARSA leaves 9 of 25 states unconverged at episode 500,000, with a maximum error of 5.53 concentrated at the state farthest from the goal. REINFORCE leaves 15 of 25 states unconverged (max error 5.71), NPG leaves 12 of 25 unconverged (max error 9.72), and PPO leaves 16 of 25 unconverged (max error 5.42).
@@ 5.4 Optimal Returns Without Optimal Value Functions
Figure 4 shows the corresponding policy grids. Q-learning’s arrows match $\pi^{*}$ everywhere by the final checkpoint. NPG and PPO show incorrect arrows concentrated at the same off-path states where value errors are largest: the agent has never needed to act at these states, so its policy there reflects the softmax initialization rather than learned behavior.
@@ 6 Reinforcement Learning for Optimal Control
Several industries now run reinforcement learning in production, not just in simulation. I focus on systems where RL operates continuously in the field, affecting millions of real decisions, such as ride-hailing dispatch at DiDi, data center cooling at Google, hotel revenue management, and financial order execution. The chapter concludes with a simulation study on the bus engine replacement problem, demonstrating that RL matches and surpasses dynamic programming on a classical benchmark.
@@ 6.1 Ride-Hailing Dispatch
Qin et al. (2021) formalized DiDi’s order dispatching as a semi-Markov decision process [^80] A semi-Markov decision process generalizes the standard MDP by allowing variable time between decisions. The discount factor $\gamma^{\tau_{k}}$ depends on the actual time elapsed $\tau_{k}$ rather than applying a fixed per-step discount. where each driver is an independent agent. The state of a driver consists of location (discretized into hexagonal zones) and time (bucketed into intervals). The action is the order assigned to the driver, with the option to remain idle. The reward is the trip fare. State transitions are determined by trip destinations, as completing an order transports the driver from origin to destination, changing their spatial state. The stochasticity arises from future demand, which determines the available actions at each state. The per-driver value function $V^{\pi}(s)$ represents the expected cumulative fare a driver can earn from state $s$ under dispatching policy $\pi$:
@@
The deployed system uses generalized policy iteration within a combinatorial optimization framework. Each dispatching window (a few seconds), the platform collects open orders and available drivers, constructs a bipartite graph, and solves a linear assignment problem. The edge weights are computed as the advantage of each driver-order match relative to the driver’s current state value:
@@
where $\hat{r}_{i}$ is the predicted fare for order $i$, $\hat{\tau}_{i}$ is the estimated trip duration, and $s_{j}^{\prime}$ is the destination state. The value function is learned via temporal-difference methods from historical trip data. DiDi’s Cerebellar Value Network (CVNet; Tang et al., 2019) uses hierarchical coarse-coding [^81] Coarse-coding represents a continuous state using overlapping receptive fields at multiple resolutions. At coarser levels, each tile covers a large region, capturing broad spatial patterns; finer levels capture local detail. The value function is the sum of features activated across all resolution levels. with a multi-resolution hexagonal grid, enabling transfer learning across cities and robustness to data sparsity in low-traffic zones.
Production deployment across more than 20 Chinese cities demonstrated consistent improvements. Table 6 summarizes the reported gains from A/B tests using time-slice rotation, where algorithms alternate control of the platform in 3-hour blocks to avoid interference effects.
@@
Han et al. (2022) reported complementary results from Lyft, where the dispatching system optimizes driver assignment and repositioning jointly. Their value decomposition architecture [^82] Value decomposition decomposes the platform’s global matching objective into individual driver value functions, each estimable via temporal-difference learning. The global optimum is recovered by optimizing the sum of individual values. assigns credit to individual driver decisions within the global objective. The production system demonstrated improvements in rider wait times and driver utilization, confirming that the RL dispatch paradigm generalizes across platforms and regulatory environments.
The percentage improvements appear modest, but at the scale of ride-hailing operations they translate to substantial aggregate effects. A 1% improvement in fulfillment rate at DiDi’s volume represents hundreds of thousands of additional completed rides per day. The central lesson from these deployments is that dispatching is not just a matching problem but a fleet positioning problem. Today’s assignments determine tomorrow’s driver distribution.
@@ 6.2 Data Center Cooling
Data centers consume enormous amounts of energy, with cooling systems accounting for 30–40% of total power usage. Google applied machine learning to autonomous HVAC control, achieving substantial efficiency gains while maintaining temperature safety constraints. The problem is a continuous control task with high-dimensional state and action spaces, strict safety requirements, and non-stationary dynamics.
Gao (2014) reported initial results from deploying neural network predictions to guide data center operations. The system used over 100 sensor readings, including temperatures at multiple locations, coolant flow rates, power draw, and external weather conditions, to predict Power Usage Effectiveness (PUE). [^83] PUE is the ratio of total facility energy to IT equipment energy. A PUE of 1.0 means all energy goes to computing; the industry average is approximately 1.6. Operators used these predictions to adjust equipment setpoints. This assisted approach achieved a 40% reduction in cooling energy and a 15% improvement in overall PUE.
Lazic et al. (2018) subsequently deployed a fully autonomous system using model-predictive control with learned dynamics. The state space consists of temperatures, flow rates, and equipment states across multiple zones. Actions are setpoints for chillers, fans, and pumps. The system learns a neural network dynamics model $\hat{f}$ from operational data:
@@
At each control interval (5 minutes), the system optimizes over a horizon of future states using the learned model, subject to temperature constraints:
@@
The system operates continuously across multiple data centers, processing millions of sensor readings per day and controlling thousands of setpoints. The learned model adapts online to seasonal variations, equipment degradation, and workload changes. The lesson is that model-based RL with explicit safety constraints is better suited to critical infrastructure than model-free approaches that require extensive exploration.
@@ 6.3 Hotel Revenue Management
Hotels face a dynamic pricing problem with perishable inventory, since rooms unsold by check-in yield zero revenue. Classical revenue management uses booking limit controls derived from dynamic programming under parametric demand assumptions. RL offers an alternative that learns pricing policies from data without specifying demand functional forms.
The problem is a finite-horizon MDP. The state $(t,I,d)$ consists of days until check-in, remaining rooms, and a demand intensity indicator. Actions are discrete price levels. Booking arrivals follow a Poisson process with intensity depending on demand conditions; each arrival books with probability decreasing in price. The objective is to maximize total revenue:
@@
The DQN agent achieves 98.7% of the DP optimal revenue while scaling to capacities where exact DP is infeasible. [^84] EMSR-b (Expected Marginal Seat Revenue, variant b) is the standard heuristic in airline and hotel revenue management. It computes booking limits by comparing the expected marginal revenue of reserving a unit for a higher fare class against selling it now. Training time is nearly constant in hotel capacity, whereas DP computation grows linearly. The learned policies exhibit interpretable patterns, with prices increasing as inventory depletes and as check-in approaches, with larger adjustments during high-demand periods.
@@ 6.4 Financial Order Execution
Optimal execution of large equity orders requires balancing execution risk (adverse price movements while holding inventory) against market impact (price deterioration from trading aggressively). The theoretical benchmark is the Almgren-Chriss framework (Almgren and Chriss, 2001), which derives optimal deterministic schedules under linear impact assumptions. Real markets exhibit state-dependent liquidity that these assumptions miss.
Nevmyvaka et al. (2006) applied tabular Q-learning to execution on real limit order book data from NASDAQ stocks (AMZN, QCOM, NVDA). The state consists of time remaining, inventory remaining, bid-ask spread, and signed volume imbalance. [^85] The limit order book records all outstanding buy and sell orders at each price level. Signed volume imbalance is the difference between buy and sell order volume near the best prices, normalized by total volume; positive imbalance typically predicts short-term price increases. Actions are trade sizes or the decision to wait. The reward is the negative implementation shortfall, [^86] Implementation shortfall measures the total cost of executing a trade relative to a benchmark price, typically the mid-quote at order arrival. It captures market impact, timing cost, and opportunity cost of unexecuted shares. the difference between execution price and a benchmark. The agent learns to condition on market microstructure signals, trading aggressively when spreads are narrow, wait when order flow predicts favorable price movement, accelerate near the deadline.
@@
The negative percentages indicate lower costs for the RL agent. Improvements of 12–19% over the Almgren-Chriss baseline are economically significant. For institutional traders executing billions of dollars annually, this represents substantial savings. Several investment banks have subsequently deployed production execution algorithms built on these ideas, using deep RL with richer state representations.
@@ 6.5 Supply Chain Inventory Management
The MDP formulation for a serial $K$-echelon system is as follows. The state $s=(I_{1},\ldots,I_{K})$ records on-hand inventory at each stage, where stage 1 faces customer demand and stage $K$ is the most upstream. The action $q\in\{0,1,\ldots,Q_{\max}\}$ is the order quantity placed at stage $K$. Demand $D_{t}\sim F_{D}$ arrives at stage 1 each period; unfilled demand is backordered. Shipments flow downstream, with stage $k$ receiving what stage $k+1$ shipped in the previous period. The per-period cost combines holding, backorder, and ordering components:
@@
Gijsbrechts et al. (2022) conducted a systematic evaluation of deep RL against classical base-stock policies across lost sales, dual sourcing, and multi-echelon configurations. The classical benchmark is the echelon base-stock policy, in which each stage $k$ maintains an echelon inventory position and orders to bring this position to a target level $S_{k}$. [^87] Formally, the echelon inventory position at stage $k$ equals on-hand inventory at $k$ plus all inventory at downstream stages $1,\ldots,k-1$ plus in-transit inventory, minus backorders. The echelon base-stock policy of Clark and Scarf (1960) is optimal for serial systems with linear costs and backorders. For single-echelon systems with backorders, the optimal base-stock level is given by the newsvendor critical fractile $S^{*}=F_{D}^{-1}(b/(b+h))$. Table 10 summarizes results from their multi-echelon experiments.

### 15 Mar 2026 v2

@@ 7.1 Setup
The environment is a $5\times 5$ deterministic gridworld with five actions (up, down, left, right, stay). Each non-terminal step incurs a penalty of 0.1-0.1; reaching the goal in the bottom-right corner yields +10+10. A small state-dependent perturbation ($\epsilon=0.001$) breaks the symmetry between equally-long paths, ensuring a unique optimal policy.113113113Without the perturbation, any shortest path from start to goal is optimal. The perturbation subtracts $\epsilon\cdot i/N^{2}$ from each step reward, where $i$ is the destination state index. Since moving down increases the state index by $N$ while moving right increases it by 1, the agent strictly prefers right over down at every interior state. The discount factor is $\gamma=0.95$ and each episode is capped at 50 steps. The optimal path requires 8 steps, yielding an optimal discounted return of 9.30.114114114Shared hyperparameters across all learning algorithms: learning rate $\alpha=0.1$ (decay 0.99995/episode) for value-based methods, $\alpha=0.01$ for policy gradient methods; $\epsilon$-greedy exploration from 1.0 to 0.01 (decay 0.9995/episode); trace decay $\lambda=0.9$; NPG actor step size $\eta=0.2$, critic step size 0.5 (both decay 0.99995/episode). DQN uses a two-layer network with 64 hidden units, replay buffer of 50,000 transitions, batch size 64, learning rate $10^{-3}$, target network update every 100 episodes, and $\epsilon$-decay 0.9999. PPO uses clip ratio 0.2, 4 optimization epochs per batch, GAE $\lambda=0.95$, learning rate $0.01$ (decay 0.99995/episode). All learning algorithms run for 500,000 episodes with seed 42.
@@ 7.2 Planning Versus Learning
VI converges in 9 iterations (0.001 seconds). On this deterministic gridworld, the iteration count matches the MDP diameter: the longest shortest path from any state to the goal is 8 steps, and each Bellman backup propagates value one step outward from the terminal state.115115115The $\gamma^{k}$ contraction bound from Section 5.1 predicts roughly $\log(\epsilon^{-1})/\log(\gamma^{-1})\approx 360$ iterations for $\epsilon=10^{-8}$, making it loose by a factor of 40 on this structured problem.
@@ 7.3 Value Function Convergence
Q-learning and Q($\lambda$) are off-policy and learn $Q^{*}$ regardless of the behavior policy. Both achieve $\max_{s}|V(s)-V^{*}(s)|<0.1$ by episode 610 and 550 respectively, with final RMSE below 0.01. Q($\lambda$) converges slightly faster due to eligibility traces propagating the terminal reward backward through the full 8-step trajectory within a single episode.
@@
SARSA, REINFORCE, NPG, and PPO are on-policy and learn the value of the exploration policy $\pi_{\epsilon}$, not $\pi^{*}$. SARSA leaves 9 of 25 states unconverged at episode 500,000, with a maximum error of 5.53 concentrated at the state farthest from the goal. REINFORCE leaves 15 of 25 states unconverged (max error 5.71), NPG leaves 12 of 25 unconverged (max error 9.72), and PPO leaves 16 of 25 unconverged (max error 5.42).
@@ 7.4 Optimal Returns Without Optimal Value Functions
Figure 14 shows the corresponding policy grids. Q-learning’s arrows match $\pi^{*}$ everywhere by the final checkpoint. NPG and PPO show incorrect arrows concentrated at the same off-path states where value errors are largest, since the agent has never needed to act at these states, so its policy there reflects the softmax initialization rather than learned behavior.
@@ 8 Reinforcement Learning for Optimal Control
A handful of organizations have deployed reinforcement learning beyond simulation, achieving measurable gains on specific large-scale problems. Each deployment required substantial domain engineering and scientific tuning; these remain exceptional cases rather than standard practice. I review the most prominent field deployments, including ride-hailing dispatch at DiDi, data center cooling at Google, hotel revenue management, and financial order execution, before concluding with a simulation study on the bus engine replacement problem.
@@ 8.1 Ride-Hailing Dispatch
Qin et al. (2021) formalized DiDi’s order dispatching as a semi-Markov decision process116116116A semi-Markov decision process generalizes the standard MDP by allowing variable time between decisions. The discount factor $\gamma^{\tau_{k}}$ depends on the actual time elapsed $\tau_{k}$ rather than applying a fixed per-step discount. where each driver is an independent agent. The state of a driver consists of location (discretized into hexagonal zones) and time (bucketed into intervals). The action is the order assigned to the driver, with the option to remain idle. The reward is the trip fare. State transitions are determined by trip destinations, as completing an order transports the driver from origin to destination, changing their spatial state. The stochasticity arises from future demand, which determines the available actions at each state. The per-driver value function $V^{\pi}(s)$ represents the expected cumulative fare a driver can earn from state $s$ under dispatching policy $\pi$:
@@
Each dispatching window (a few seconds), the platform collects open orders and available drivers, constructs a bipartite graph, and solves a linear assignment problem. The edge weights are computed as the advantage of each driver-order match relative to the driver’s current state value.
@@
where $\hat{r}_{i}$ is the predicted fare for order $i$, $\hat{\tau}_{i}$ is the estimated trip duration, and $s_{j}^{\prime}$ is the destination state. The value function is learned via temporal-difference methods from historical trip data. DiDi’s Cerebellar Value Network (CVNet; Tang et al., 2019) uses hierarchical coarse-coding117117117Coarse-coding represents a value function as a weighted sum over overlapping rectangular or hexagonal tiles that partition the state space (Sutton and Barto, 2018); each state activates the tiles that contain it, and a hierarchical variant stacks tiles at multiple resolutions to allow both coarse and fine generalization simultaneously. with a multi-resolution hexagonal grid, enabling transfer learning118118118Transfer learning initializes a model for a new task (a new city) with parameters trained on related tasks (existing cities), on the assumption that learned representations of traffic and demand patterns generalize across contexts and require less data to reach good performance in the new setting. across cities and robustness to data sparsity in low-traffic zones.
Production deployment across more than 20 Chinese cities demonstrated modest but consistent improvements of 0.5–2% on key metrics, gains that required the full CVNet infrastructure (hierarchical coarse-coding, multi-resolution grids, transfer learning across cities) to achieve. Table 11 summarizes the reported gains from A/B tests using time-slice rotation, where algorithms alternate control of the platform in 3-hour blocks to avoid interference effects.
@@
Han et al. (2022) reported complementary results from Lyft, where the dispatching system optimizes driver assignment and repositioning jointly. Their value decomposition architecture119119119Value decomposition decomposes the platform’s global matching objective into individual driver value functions, each estimable via temporal-difference learning. The global optimum is recovered by optimizing the sum of individual values. assigns credit to individual driver decisions within the global objective. The production system demonstrated improvements in rider wait times and driver utilization, providing a second data point suggesting that similar approaches may transfer across platforms.
At DiDi’s volume, even 1% gains represent hundreds of thousands of additional completed rides per day, because dispatching is fundamentally a fleet positioning problem: today’s assignments determine tomorrow’s driver distribution.
@@ 8.2 Data Center Cooling
Data centers consume enormous amounts of energy, with cooling systems accounting for 30–40% of total power usage. Google applied machine learning to autonomous HVAC control, achieving substantial efficiency gains while maintaining temperature safety constraints.
Gao (2014) reported initial results from deploying neural network predictions to guide data center operations. The system used over 100 sensor readings, including temperatures at multiple locations, coolant flow rates, power draw, and external weather conditions, to predict Power Usage Effectiveness (PUE).120120120PUE is the ratio of total facility energy to IT equipment energy. A PUE of 1.0 means all energy goes to computing; the industry average is approximately 1.6. Operators used these predictions to adjust equipment setpoints. This assisted approach achieved a 40% reduction in cooling energy and a 15% improvement in overall PUE.
Lazic et al. (2018) subsequently deployed a fully autonomous system using model-predictive control121121121Model-predictive control (MPC) solves a finite-horizon optimal control problem at each step using a predictive model of system dynamics, applies the first action, then re-solves at the next step with the updated state; the receding-horizon structure makes it robust to model error because the plan is continually revised. with learned dynamics. The state space consists of temperatures, flow rates, and equipment states across multiple zones. Actions are setpoints for chillers, fans, and pumps. The system learns a neural network dynamics model $\hat{f}$ from operational data:
@@
At each control interval (5 minutes), the system optimizes over a horizon of future states using the learned model, subject to temperature constraints.
@@
The system operates continuously across multiple data centers, processing millions of sensor readings per day and controlling thousands of setpoints. Importantly, this is model-predictive control with learned dynamics models and persistent human oversight, not autonomous model-free RL; the learned model adapts online to seasonal variations, equipment degradation, and workload changes, but operators retain override authority at all times.
@@ 8.3 Hotel Revenue Management
Hotels face a dynamic pricing problem with perishable inventory, since rooms unsold by check-in yield zero revenue. The state $(t,I,d)$ consists of days until check-in, remaining rooms, and a demand intensity indicator. Actions are discrete price levels. Booking arrivals follow a Poisson process with intensity depending on demand conditions; each arrival books with probability decreasing in price. The Bellman equation is:
@@
The DQN agent achieves 98.7% of the DP optimal revenue while scaling to capacities where exact DP is infeasible.122122122EMSR-b (Expected Marginal Seat Revenue, variant b) is the standard heuristic in airline and hotel revenue management. It computes booking limits by comparing the expected marginal revenue of reserving a unit for a higher fare class against selling it now. Training time is nearly constant in hotel capacity, whereas DP computation grows linearly. The learned policies exhibit interpretable patterns, with prices increasing as inventory depletes and as check-in approaches, with larger adjustments during high-demand periods.
@@ 8.4 Financial Order Execution
The theoretical benchmark for optimal execution is the Almgren-Chriss framework (Almgren and Chriss, 2001), which derives optimal deterministic schedules under linear impact assumptions. A trader liquidating $Q$ shares over $T$ periods faces a tradeoff between timing risk and market impact. With risk-aversion parameter $\lambda$, price volatility $\sigma$, and temporary impact coefficient $\eta$, the optimal remaining inventory at time $t$ follows the hyperbolic sine schedule:
$x^{*}(t)=Q\cdot\frac{\sinh\!\bigl(\kappa(T-t)\bigr)}{\sinh(\kappa T)},\quad\kappa=\sqrt{\frac{\lambda\sigma^{2}}{\eta}}.$ (64)
This deterministic trajectory is the benchmark any adaptive method must beat. It prescribes front-loading or back-loading depending on the risk-impact balance, but cannot respond to realized order flow or spread dynamics.
Nevmyvaka et al. (2006) applied tabular Q-learning to execution on real limit order book data from NASDAQ stocks.123123123The dataset covers 500 trading days of millisecond-level limit order book snapshots for AMZN, QCOM, and NVDA. The state space contains approximately 10,000 states; the horizon is $T=60$ seconds with discount $\gamma=1$. The state at time $t$ is $s_{t}=(t,q_{t},\psi_{t},\Delta_{t})$, where $t\in\{1,\ldots,T\}$ is time remaining, $q_{t}\in\{0,1,\ldots,Q\}$ is inventory remaining, $\psi_{t}$ is the discretized bid-ask spread, and $\Delta_{t}$ is the discretized signed volume imbalance.124124124Signed volume imbalance is the difference between buy and sell order volume near the best prices, normalized by total volume; positive imbalance typically predicts short-term price increases. Actions $a_{t}\in\{0,\delta,2\delta,\ldots,q_{t}\}$ specify shares to execute in the current interval. The reward is the negative per-period slippage contribution:
$r_{t}=-a_{t}\bigl(p_{t}^{\text{exec}}-p_{0}\bigr),$ (65)
where $p_{0}$ is the mid-quote at order arrival and $p_{t}^{\text{exec}}$ is the average execution price. Total implementation shortfall is $IS=-\sum_{t=1}^{T}r_{t}/Q$.125125125Implementation shortfall measures the total cost of executing a trade relative to the benchmark mid-quote at arrival. It captures market impact, timing cost, and opportunity cost of unexecuted shares. Because the objective is cost minimization, the Q-learning update uses $\min$ over next-period actions:
$Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\alpha\Bigl[r_{t}+\gamma\min_{a^{\prime}}Q(s_{t+1},a^{\prime})-Q(s_{t},a_{t})\Bigr].$ (66)
The agent learns to condition on market microstructure signals: trading aggressively when spreads are narrow, waiting when order flow predicts favorable price movement, and accelerating near the deadline.
@@
The RL agent reduces execution costs by 12–19% over Almgren-Chriss. These results informed subsequent work on adaptive execution, though independently verified production deployments remain scarce in the public literature.
@@ 8.5 Supply Chain Inventory Management
The state $s=(I_{1},\ldots,I_{K})$ records on-hand inventory at each stage, where stage 1 faces customer demand and stage $K$ is the most upstream. The action $q\in\{0,1,\ldots,Q_{\max}\}$ is the order quantity placed at stage $K$. Demand $D_{t}\sim F_{D}$ arrives at stage 1 each period; unfilled demand is backordered. Shipments flow downstream, with stage $k$ receiving what stage $k+1$ shipped in the previous period. The per-period cost combines holding, backorder, and ordering components.
@@
Gijsbrechts et al. (2022) conducted a systematic evaluation of deep RL against classical base-stock policies across lost sales, dual sourcing, and multi-echelon configurations. The classical benchmark is the echelon base-stock policy, in which each stage $k$ maintains an echelon inventory position and orders to bring this position to a target level $S_{k}$.126126126Formally, the echelon inventory position at stage $k$ equals on-hand inventory at $k$ plus all inventory at downstream stages $1,\ldots,k-1$ plus in-transit inventory, minus backorders. The echelon base-stock policy of Clark and Scarf (1960) is optimal for serial systems with linear costs and backorders. For single-echelon systems with backorders, the optimal base-stock level is given by the newsvendor critical fractile $S^{*}=F_{D}^{-1}(b/(b+h))$. Table 15 summarizes results from their multi-echelon experiments.